# Exploring the cosmological degeneracy between decaying dark matter model and viscous $\Lambda$CDM

### v1

@@ Abstract
In the context of an homogeneous and isotropic universe, we consider the degeneracy condition at background level between two scenarios in which processes out of equilibrium are possible to explore their cosmological implications; this consideration allows us to deviate from the perfect fluid description and in this case bulk viscosity represents a viable candidate to describe entirely such effects. The cosmological model describing an unstable dark matter sector is mapped into a slight modification of the $\Lambda$CDM model characterized by a viscous dark matter sector; under this consideration our description does not depend on a specific formulation of viscous effects and these can be fully reconstructed and characterized by the parameter that determines the decay ratio of dark matter. However, in this new scenario the cosmic expansion is influenced by the viscous pressure and the dark energy sector given by the cosmological constant is translated into a dynamical one. As consequence of our formulation, the test against observations of the model indicates consistency with quintessence dark energy but the crossing of the phantom divide can be accessible.
@@ I Introduction
It is well known that the cosmological concordance model, or simply $\Lambda$CDM, is based on the existence of a non-relativistic fluid known as cold dark matter and a cosmological constant (CC) which is interpreted as vacuum energy and represents the dark energy sector. In this scenario dark matter dominates the growth of structures and deceleration at early stages of cosmological evolution, while the CC drives the current accelerated expansion. The amount of dark matter at the recombination epoch has been inferred from the Cosmic Microwave Background (CMB) observations planck1518 and its comoving density is constant throughout the cosmic evolution together with equation of state, $\omega_{m}=0$; besides, the linearized density and velocity perturbations satisfy the continuity and pressureless Euler equations. However, as attractive as the simplicity of the $\Lambda$CDM model may be, the high-definition astrophysical data indicates that we require to look beyond this model in order to describe the observable universe more accurately riess2024 .
@@
Regarding the aforementioned necessity and leaving the dark energy sector aside for the moment, the most simple assumption to explore is the consideration of deviations from the usual description mentioned above for dark matter. As first example we can mention the generalized dark matter model given by Hu hu , in which $\omega_{m}\neq 0$ through the cosmic history, this proposal was extensively studied and extended to the context of interacting dark sector in hu2 ; hu3 . The results reported in Ref. hu3 indicated that the dark matter abundance is strictly positive and $\omega_{m}$ is consistent with zero. But, an interesting characteristic of this model is that the non-linearities at small scales can alter the cosmological background and large scale linear perturbations, these effects are characterized by the creation of an effective pressure and viscosity in the fluid, the physical interpretation is that viscosity will act to damp out the velocity fluctuations. The emergence of viscosity (and effective pressure) in an initially pressureless perfect fluid was also confirmed in the context of an effective field theory of large scale structure large , i.e., the long-wavelength universe behaves as a viscous fluid coupled to gravity.
@@
Another interesting possibility to consider is the relaxation of the comovingly constant dark matter density condition. A captivating approach is the decay of a fraction of dark matter into relativistic particles of the dark sector, i.e., dark radiation. This is the most simple scenario and seems to resolve several discrepancies of standard cold dark matter, see for instance takahashi . On the other hand, such a conversion has been subject of exhaustive investigation since alleviates some of the tensions that plague the $\Lambda$CDM model, see Refs. lesgourgues ; das1 ; das2 ; bringmann ; vattis and references therein. Additionally, the conversion of dark matter into dark radiation can be also found in the merging of primordial black holes emitting gravitational waves bh , this can promote the use of the recent results released by the LIGO collaboration ligo in order to be implemented in this kind of cosmological scenario. The nature of dark radiation is not completely known so far. However, some studies allow dark radiation to be comprised by sterile neutrinos as well as active neutrinos with a non-thermal distribution, in such case neutrinos are expected to behave as relativistic particles with effective sound speed and viscosity parameter, see Ref. lesgourgues2 ; the existence of bulk viscosity in the universe was demonstrated long time ago using a gas of neutrinos neutribulk . Notice that in both scenarios mentioned here for modified dark matter, the emergence of a viscous contribution in the cosmic fluid is a common factor.
@@
In this sense, in order to study the cosmological viability of both approaches to describe the current accelerated expansion; in this work we consider a mapping at background level between a specific cosmological model in which dissipative processes are present due to the slow decay of dark matter into dark radiation with a dark energy sector characterized by a constant parameter state, $\omega_{\mathrm{de}}$ wcdm and a slightly modified version of the $\Lambda$CDM model, in this case the modification is given by the consideration of a viscous dark matter sector; due to the contribution of the viscous pressure, $\Pi$, the condition $\omega_{m}\neq 0$ is covered. As a result of this degeneracy condition, the viscous effects considered in the $\Lambda$CDM model can be characterized by a single parameter, $\alpha$, which in turn determines the ratio of decay for dark matter and at the same time the CC becomes dynamical. This latter situation can be useful to circumvent the well-known CC problem: the great discrepancy found between the values of $\Lambda$ (interpreted as vacuum energy) obtained from quantum field theory and observations indicates that the nature of dark energy can not be entirely comprised by the CC lambda . The conversion from the constant behavior to a dynamical one for the dark energy sector is obtained naturally in this formulation, but is the most simple phenomenological way to extend and explain the nature of dark energy and has been widely explored in the literature by considering parametrizations of the form, $\Lambda=\Lambda(a)$ and/or $\Lambda=\Lambda(H)$, being $a$ the scale factor and $H$ the Hubble parameter, see for instance Refs. given in vcc . Additionally, this procedure allows to reconstruct the form of the bulk viscosity without need of solving the constitutive equations that characterize the viscous effects, in general these equations are hard to solve in this kind of scenario is ; nonlinear . The presence of the term $\Pi$ lowers the total pressure of the fluid and in consequence contributes to the accelerated cosmic expansion.
@@
We highlight the fact that both models to be considered in this work have in common that their processes occur out of equilibrium and the mapping can be done given the degeneracy of the expansion of the universe at background level following the methodology given in Ref. velten for an homogeneous and isotropic universe.
@@
The structure of the article is as follows: in the next Section the generalities of the effective cosmological models to be used are described, as well as the properties of the thermodynamics of irreversible processes involved, we comment the similarities at background dynamics level between bulk viscosity and the matter creation scenario. We specify the main properties of the decaying dark matter model and the modified $\Lambda$CDM model. In Section III a Bayesian analysis is carried out in order to test the model against observational data. In this analysis, the parameters of the model are constrained and their uncertainties are estimated. Finally in section IV we give some final comments for our work.
@@ II Cosmological models
This section is devoted to describe the general aspects of the cosmological models to consider in our analysis for a Friedmann-Lemaitre-Robertson-Walker (FLRW) space-time with null spatial curvature, $18\pi G=c=$ 1 units will be used.
@@ II.1 Some aspects of irreversible thermodynamics
In general, the evolution of the universe is described assuming that matter corresponds to a perfect fluid. However, by considering a fluid of this type, some physical features of matter are left out. Such features are relevant if dissipative processes occur. Consequently, under the typical prescription for matter, cosmic evolution results to be a reversible process. Within such standard scenario, physical states of matter are described by local equilibrium variables even when processes for which the non-equilibrium condition is expected are undergoing, as happens for the expanding fluid case. Clearly, something is missing in its description from the thermodynamics point of view. In order to generalize the description of the cosmic fluid, in this work we consider deviations from local equilibrium variables. A well known example of this kind of schema is to introduce dissipative effects. For a dissipative fluid, the energy density coincides with the local equilibrium value (this is also valid for other thermodynamics scalars), but the pressure deviates from the local pressure value as follows maartens
@@
where the dot denotes derivatives w.r.t. proper time and $H$ is the Hubble parameter. Notice that the above equation can be written as an inhomogeneous equation, being $\Pi$ the source/sink of energy density. Such role for $\Pi$ has an important meaning at thermodynamics level since it also can be though as an entropy catalyst, leading to a non adiabatic cosmological expansion, as we will see below. Therefore this description for the fluid does not satisfy the equilibrium condition. In this case the Friedmann equations are modified in the following form
@@
An interesting fact about the set of equations describing the cosmic evolution of dissipative fluids is their equivalence with those when matter-creation effects arise within homogeneous spacetimes, as discussed in Ref. zimdahl from the kinetic theory perspective, see also Refs. given in equiv . In this case the particle production is described by the particle creation pressure, $p_{c}$, which in turn is characterized by a $\Gamma$ term. Such pressure plays the role of $\Pi$ in equations (1), (2) and (4), i.e., we now write $p_{\mathrm{eff}}=p+p_{c}$ for the effective pressure and from this expression we can observe, $p_{c}\leftrightarrow\Pi$. From now on we will characterize the deviations from equilibrium pressure simply as $\Pi$ for simplicity in the notation. For the matter creation scenario we must also consider that the particle number is non conserved
@@ II.2 Decaying dark matter plus dark energy
where $\rho_{\mathrm{m}}$ denotes the dark matter energy density and $\rho_{\mathrm{dr}}$ denotes the dark radiation generated by the dark matter decay. In this case we have considered a barotropic EoS, $p_{\mathrm{i}}=\omega_{\mathrm{i}}\rho_{\mathrm{i}}$, where the subscript $i$ denotes different components, notice that we are restricting ourselves to cold dark matter case, i.e., $\omega_{\mathrm{m}}=0$. Besides $Q$ denotes the decaying rate for dark matter. In general the form of the decay rate is given by, see for instance das1 ; das2
@@
where $\Delta=\alpha H$, being $\alpha$ a positive constant. If $\alpha\sim 1$ then mean lifetime of dark matter is about a Hubble time at different epochs. For the dark energy sector we consider the following continuity equation wcdm
@@
where $\omega_{\mathrm{de}}$ is a constant parameter of state associated to dark energy which is left as a free parameter in order to track effects on its values due to the occurrence of a irreversible process such as dark matter decay. As can be seen from the previous equations, the total energy density is conserved while those related to specific species do not. Adopting the standard relationship between the scale factor $a$ and the red-shift $z$ given as $1+z=a^{-1}$, we can perform a change of variable in the system of Eqs. (9) and (10), then by considering the previous expressions for the decay rate we can obtain the following analytic solutions for the energy densities for both components das1 ; das2
@@
being $\rho_{\mathrm{i,0}}$ the value of energy density at present time, $\rho_{\mathrm{i}}(z=0)=\rho_{\mathrm{i,0}}$. Then, for this cosmological model the Friedmann constraint reads $3H^{2}(z)=(\rho_{\mathrm{m}}(z)+\rho_{\mathrm{dr}}(z)+\rho_{\mathrm{de}}(z))$ and defining the normalized Hubble parameter, $E(z):=H(z)/H_{0}$ with $H_{0}$ being the Hubble constant give as $H(z=0):=H_{0}$, we can write
@@ II.3 Modified $\Lambda$CDM model
Following the line of reasoning of Ref. velten , we now proceed to explore the degeneracy problem at background level between the cosmological model (16) and a $\Lambda$CDM model with a modified dark matter sector. This means that if both cosmological models give rise to the same expansion for the universe then we must have, $E(z)=E_{\Lambda}(z)$. Consequently both parameter spaces can be mapped to each other. The modification made to the dark matter sector of the $\Lambda$CDM model is given by the consideration of an effective non-perfect pressure $p_{\mathrm{eff}}$ instead of the local equilibrium pressure, then in our analysis the $\Pi$ term could be associated to an effective viscous pressure or a matter creation pressure. We first focus on the viscous scenario. Under the considerations commented above the pressure of matter sector takes the form
@@
we have used Eq. (4) to write the viscous pressure as, $\Pi=\rho_{\Lambda}-\frac{1}{3}\rho_{r}-3H^{2}-2\dot{H}^{2}$. From the degeneracy condition $E(z)=E_{\Lambda}(z)$, one gets the following expression for the viscous dark matter
@@
Therefore if we insert the result given above in Eq. (19) one gets the following expression
@@
If we consider the conditions given in velten we have, $\Omega_{\mathrm{r,0}}=\bar{\Omega}_{\mathrm{r,0}}$ since the density of radiation fluid is obtained from the temperature of the CMB, in other words, this parameter is not model dependent. Besides, if the correct amount of matter is established correctly from observations then $\Omega_{\mathrm{m,0}}=\Omega_{\mathrm{vm,0}}$, which in turn implies that dark energy density parameter will be also well determined for any model, then $\Omega_{\mathrm{de,0}}=\Omega_{\Lambda}$. Therefore Eq. (19) takes the form
@@
where the r.h.s. of Eq. (19) is replaced by the l.h.s. of Eq. (21). Notice that the viscous matter sector can be written in terms of the parameters of the cosmological model (16). It is worthy to mention that the solution for $\Omega_{\mathrm{vm}}(z)$ does not depend on a specific election of $\Pi$, then this solution would be valid to characterize viscous effects obtained from the full theory of the causal Israel-Stewart formalism is and even from the non linear extension formalism given in nonlinear , which is more adequate to describe an expanding fluid. Since matter creation effects are uniquely characterized by deviations from the equilibrium pressure, then Eq. (22) can also describe the created matter scheme, $\Omega_{\mathrm{vm}}(z)\longrightarrow\Omega_{\mathrm{cm}}(z)$, both scenarios will provide a solution for the matter sector in terms of the parameters associated to the decaying dark matter model plus dark energy $(\alpha,\omega_{\mathrm{de}})$ given in Eq. (16). In the context of matter creation scenario, the Ansatz for the $\Gamma$ term is necessary in order to determine the form of the creation pressure. However, as can be seen in our description, we can dispense of the explicit form of this term. See for instance Ref. gamma for examples of $\Gamma$.
@@
We would like to highlight the fact that under the inclusion of viscous effects and the consideration of the degeneracy condition at background level between both models, the expansion of the universe within the viscous $\Lambda$CDM model (23) behaves exactly as in the decaying dark matter model given in (16). However, the physical interpretation is distinct for each case; as can be seen, the modified $\Lambda$CDM model has a matter sector with a viscous pressure that can be modeled by means of the decaying dark matter model parameters (as we will see below) and the nature of the CC changes into a dynamical behavior, $\Lambda\rightarrow\Lambda(z)$, characterized by the parameter state $\omega_{\mathrm{de}}$.
In Fig. (1) we show the normalized Hubble parameter (23) and compare it to $\Lambda$CDM model for the recent evolution of the universe. As can be seen, in the past the cases are practically indistinguishable for a small value of $\alpha$. However, as the value of $\alpha$ increases we can observe deviations from $\Lambda$CDM at the past for $z>1$. Some other differences with respect to $\Lambda$CDM arise in the far future ($z=-1$), whereas the Hubble parameter for $\Lambda$CDM tends to the constant value given by, $\sqrt{\Omega_{\Lambda}}$, in our model it vanishes for quintessence and diverges for phantom dark energy, therefore in this phantom scenario the singularity is kicked away, this is known as little rip scenario lr , i.e., the model does not exhibit a true (big rip) singularity at a finite cosmic time, the singularity is only allowed until infinite time has elapsed. See for instance lr2 , where is discussed that viscous effects can naturally lead to a little rip cosmic evolution avoiding all problems of a big rip singularity, see also lr3 .
It is worthy to mention that for quintessence dark energy all values of $\alpha$ considered tend to increase the value of the normalized Hubble parameter w.r.t. the $\Lambda$CDM model in the low red-shift region, $10<z<$ 1, this is an interesting characteristic since such behavior is known to alleviate some tensions of $\Lambda$CDM, see for instance Ref. vagnozzi where this kind of condition was studied for different dark energy models.
Figure 1: Behavior of the parameter $E_{\Lambda}(z)$. For comparison purposes we use the values $\Omega_{\mathrm{r,0}}=0$, $\Omega_{\mathrm{m,0}}=0.3$, $\Omega_{\mathrm{de,0}}=0.7$. We focus on the recent evolution of the universe, $z\approx 0.6$, since for that redshift value the transition decelerated-accelerated for the cosmic expansion is expected to take place moresco .

### v2

@@ Abstract
In the context of a homogeneous and isotropic universe, we consider the degeneracy condition at the background level between two scenarios in which processes out of equilibrium are possible; this consideration allows us to deviate from the perfect fluid description, and in this case bulk viscosity represents a viable candidate to describe entirely such effects. The cosmological model describing an unstable dark matter sector is mapped into a slight modification of the $\Lambda$CDM model characterized by a viscous dark matter sector; under this consideration our description does not depend on a specific formulation of viscous effects and these can be fully reconstructed and characterized by the parameter that determines the decay ratio of dark matter. However, in this new scenario the cosmic expansion is influenced by the viscous pressure and the dark energy sector given by the cosmological constant is translated into a dynamical one. As a consequence of our formulation, the test against observations of the model indicates consistency with quintessence dark energy, but the crossing of the phantom divide can be accessible.
@@ I Introduction
It is well known that the cosmological concordance model, or simply $\Lambda$CDM, is based on the existence of a non-relativistic fluid known as cold dark matter and a cosmological constant (CC) which is interpreted as vacuum energy and represents the dark energy sector. In this scenario dark matter dominates the growth of structures and deceleration at early stages of cosmological evolution, while the CC drives the current accelerated expansion. The amount of dark matter at the recombination epoch has been inferred from the Cosmic Microwave Background (CMB) observations planck1518 and its comoving density is constant throughout the cosmic evolution together with the equation of state $\omega_{m}=0$; besides, the linearized density and velocity perturbations satisfy the continuity and pressureless Euler equations. However, as attractive as the simplicity of the $\Lambda$CDM model may be, the high-definition astrophysical data indicate that we require to look beyond this model in order to describe the observable universe more accurately riess2024 .
@@
Regarding the aforementioned necessity and leaving the dark energy sector aside for the moment, the most simple assumption to explore is the consideration of deviations from the usual description mentioned above for dark matter. As first example we can mention the generalized dark matter model given by Hu hu , in which $\omega_{m}\neq 0$ through the cosmic history, this proposal was extensively studied and extended to the context of interacting dark sector in hu2 ; hu3 . The results reported in Ref. hu3 indicated that the dark matter abundance is strictly positive and $\omega_{m}$ is consistent with zero. But an interesting characteristic of this model is that the non-linearities at small scales can alter the cosmological background and large scale linear perturbations, these effects are characterized by the creation of an effective pressure and viscosity in the fluid, the physical interpretation is that viscosity will act to damp out the velocity fluctuations. The emergence of viscosity (and effective pressure) in an initially pressureless perfect fluid was also confirmed in the context of an effective field theory of large-scale structure large , i.e., the long-wavelength universe behaves as a viscous fluid coupled to gravity.
@@
Another interesting possibility to consider is the relaxation of the comovingly constant dark matter density condition. A captivating approach is the decay of a fraction of dark matter into relativistic particles of the dark sector, i.e., dark radiation. This is the most simple scenario and seems to resolve several discrepancies of standard cold dark matter; see, for instance, takahashi . On the other hand, such a conversion has been the subject of exhaustive investigation since alleviates some of the tensions that plague the $\Lambda$CDM model; see Refs. lesgourgues ; das1 ; das2 ; bringmann ; vattis and references therein. Additionally, the conversion of dark matter into dark radiation can also be found in the merging of primordial black holes emitting gravitational waves bh , this can promote the use of the recent results released by the LIGO Collaboration ligo in order to be implemented in this kind of cosmological scenario. The nature of dark radiation has not been fully understood so far. However, some studies allow dark radiation to be comprised by sterile neutrinos as well as active neutrinos with a non-thermal distribution, in such a case neutrinos are expected to behave as relativistic particles with effective sound speed and viscosity parameter, see Ref. lesgourgues2 ; the existence of bulk viscosity in the universe was demonstrated long ago using a gas of neutrinos neutribulk . Notice that in both scenarios mentioned here for modified dark matter, the emergence of a viscous contribution in the cosmic fluid is a common factor.
@@
In this sense, in order to study the cosmological viability of both approaches to describe the current accelerated expansion; in this work we consider a mapping at the background level between a specific cosmological model in which dissipative processes are present due to the slow decay of dark matter into dark radiation with a dark energy sector characterized by a constant parameter state, $\omega_{\mathrm{de}}$ wcdm and a slightly modified version of the $\Lambda$CDM model, in this case the modification is given by considering a viscous dark matter sector; due to the contribution of the viscous pressure, $\Pi$, the condition $\omega_{m}\neq 0$ is covered. As a result of this degeneracy condition, the viscous effects considered in the $\Lambda$CDM model can be characterized by a single parameter, $\alpha$, which in turn determines the decay ratio for dark matter and at the same time the CC becomes dynamical. This latter situation can be useful to circumvent the well-known CC problem: the great discrepancy found between the values of $\Lambda$ (interpreted as vacuum energy) obtained from quantum field theory and observations indicates that the nature of dark energy can not be entirely comprised by the CC lambda . The conversion from the constant behavior to a dynamical one for the dark energy sector is obtained naturally in this formulation, but is the most simple phenomenological way to extend and explain the nature of dark energy and has been widely explored in the literature by considering parameterizations of the form $\Lambda=\Lambda(a)$ and/or $\Lambda=\Lambda(H)$, being $a$ the scale factor and $H$ the Hubble parameter; see, for instance Refs. given in vcc . Additionally, this procedure allows one to reconstruct the form of the bulk viscosity without the need of solving the constitutive equations that characterize the viscous effects, in general, these equations are hard to solve in this kind of scenario is ; nonlinear . The presence of the term $\Pi$ reduces the total pressure of the fluid and, consequently, contributes to the accelerated cosmic expansion.
@@
We highlight the fact that both models to be considered in this work have in common that their processes occur out of equilibrium, and the mapping can be done given the degeneracy of the expansion of the universe at background level following the methodology given in Ref. velten for an homogeneous and isotropic universe.
@@
The structure of the article is as follows: in the next Section the generalities of the effective cosmological models to be used are described, as well as the properties of the thermodynamics of irreversible processes involved, we comment the similarities at background dynamics level between bulk viscosity and the matter creation scenario. We specify the main properties of the decaying dark matter model and the modified $\Lambda$CDM model. In Section III a Bayesian analysis is performed in order to test the model against the observational data. In this analysis, the parameters of the model are constrained and their uncertainties are estimated. Finally in section IV we give some final comments for our work.
@@ II Cosmological models
This section is devoted to describing the general aspects of the cosmological models to consider in our analysis for a Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime with null spatial curvature; $18\pi G=c=$ 1 units will be used.
@@ II.1 Some aspects of irreversible thermodynamics
In general, the evolution of the universe is described by assuming that matter corresponds to a perfect fluid. However, by considering a fluid of this type, some physical features of matter are left out. Such features are relevant if dissipative processes occur. Consequently, under the typical prescription for matter, cosmic evolution turns out to be a reversible process. Within such a standard scenario, physical states of matter are described by local equilibrium variables even when processes for which the non-equilibrium condition is expected, as happens for the expanding fluid case. Clearly, something is missing from its description from the thermodynamic point of view. In order to generalize the description of the cosmic fluid, in this work we consider deviations from local equilibrium variables. A well known example of this kind of schema is to introduce dissipative effects. The evolution of the universe contains several dissipative processes, we mention some: the spontaneous symmetry breaking when gauge bosons acquire mass, the decay of the primordial scalar field into particles, i.e., the reheating stage at the end of inflation, the decoupling of neutrinos from cosmic plasma and the decoupling of photons from matter during the recombination era. Within the astrophysical context the gravitational collapse of local inhomogeneities to form galactic structures is an example in which dissipation is involved. Additionally, the collapse of a radiating star to a neutron star o black hole is another example, since heat flow and viscosity are present. Finally, the accretion of matter around a neutron star or black hole and binary neutron star merger are also examples of dissipative processes; these latter examples provide a good environment to explore the properties of matter at extreme densities and temperatures maartens ; noronha .
For a dissipative fluid, the energy density coincides with the local equilibrium value (this is also valid for other thermodynamics scalars), but the pressure deviates from the local pressure value as follows maartens
@@
where the dot denotes derivatives w.r.t. proper time and $H$ is the Hubble parameter. Notice that the above equation can be written as an inhomogeneous equation, being $\Pi$ the source/sink of energy density. Such role for $\Pi$ has an important meaning at thermodynamics level since it also can be though as an entropy catalyst, leading to a non adiabatic cosmological expansion, as we will see below. This latter characteristic represents an interesting cosmological scenario since non-equilibrium phenomena are also present in other cosmological models. For instance, in Ref. entropy it was found that the growth of entropy associated to the causal horizon of our universe induces an acceleration that is indistinguishable from the one obtained in the $\Lambda$CDM model for an homogeneous and isotropic cosmology. Based on the aforementioned discussion, our description for the fluid does not satisfy the equilibrium condition. In this case the Friedmann equations are modified in the following form
@@
It is worthy to mention that the set given by the equations (2), (3) and (4), resembles the background dynamics associated to more general cosmological scenarios as unimodular gravity. Such scheme allows diffusion processes which induce deviations from the equilibrium condition due to the non-conservation of the energy-momentum tensor, see Ref. nucamendi for an interesting review of the model and its role alleviating the $H_{0}$ tension. An interesting fact about the set of equations describing the cosmic evolution of dissipative fluids is their equivalence with those when matter-creation effects arise within homogeneous spacetimes, as discussed in Ref. zimdahl from the kinetic theory perspective, see also Refs. given in equiv . In this case the particle production is described by the particle creation pressure, $p_{c}$, which in turn is characterized by a $\Gamma$ term. Such pressure plays the role of $\Pi$ in equations (1), (2) and (4), i.e., we now write $p_{\mathrm{eff}}=p+p_{c}$ for the effective pressure and from this expression we can observe, $p_{c}\leftrightarrow\Pi$. From now on we will characterize the deviations from equilibrium pressure simply as $\Pi$ for simplicity in the notation. For the matter creation scenario we must also consider that the particle number is non conserved
@@ II.2 Decaying dark matter plus dark energy
where $\rho_{\mathrm{m}}$ denotes the dark matter energy density and $\rho_{\mathrm{dr}}$ denotes the dark radiation generated by the dark matter decay. In this case we have considered a barotropic EoS, $p_{\mathrm{i}}=\omega_{\mathrm{i}}\rho_{\mathrm{i}}$, where the subscript $i$ denotes different components, notice that we are restricting ourselves to the cold dark matter case, i.e. $\omega_{\mathrm{m}}=0$. According to Eqs. (9) and (10), this kind of model is related to the scenario of an interacting dark sector where the $Q$-term is a function that mediates the exchange of energy between dark species with the direction of the energy flow defined by the sign of $Q$. The interacting scheme represents a generalized cosmological model since the entropy production is allowed due to the presence of the $Q$-term; this differs from the adiabatic cosmic expansion obtained from the cosmological standard model; the interacting model also gives rise to a robust thermodynamic description of the cosmic fluids since some variable temperatures can be associated to the dark sector. For interesting reviews on this topic, see Refs. given in pavon . In this case, $Q$ denotes the decay rate of dark matter. In general the form of the decay rate is given by, see for instance das1 ; das2
@@
where $\Delta=\alpha H$, being $\alpha$ a positive constant. If $\alpha\sim 1$, then the mean lifetime of dark matter is about a Hubble time at different epochs. For the dark energy sector we consider the following continuity equation wcdm
@@
where $\omega_{\mathrm{de}}$ is a constant parameter of state associated to dark energy which is left as a free parameter in order to track effects on its values due to the occurrence of a irreversible process such as dark matter decay. As can be seen from the previous equations, the total energy density is conserved, while those related to specific species do not. Adopting the standard relationship between the scale factor $a$ and the red-shift $z$ given as $1+z=a^{-1}$, we can perform a change of variable in the system of equations (9) and (10), then by considering the previous expressions for the decay rate we can obtain the following analytic solutions for the energy densities for both components das1 ; das2
@@
being $\rho_{\mathrm{i,0}}$ the value of energy density at present time, $\rho_{\mathrm{i}}(z=0)=\rho_{\mathrm{i,0}}$. Then, for this cosmological model the Friedmann constraint reads $3H^{2}(z)=(\rho_{\mathrm{m}}(z)+\rho_{\mathrm{dr}}(z)+\rho_{\mathrm{de}}(z))$ and defining the normalized Hubble parameter, $E(z):=H(z)/H_{0}$ with $H_{0}$ being the Hubble constant given as $H(z=0):=H_{0}$, we can write
@@ II.3 Modified $\Lambda$CDM model
Following the line of reasoning of Ref. velten , we now proceed to explore the degeneracy problem at the background level between the cosmological model (16) and a $\Lambda$CDM model with a modified dark matter sector. This means that if both cosmological models give rise to the same expansion for the universe, then we must have $E(z)=E_{\Lambda}(z)$. Consequently, both parameter spaces can be mapped to each other. The modification made to the dark matter sector of the $\Lambda$CDM model is given by the consideration of an effective non-perfect pressure $p_{\mathrm{eff}}$ instead of the local equilibrium pressure, and then in our analysis the $\Pi$ term could be associated to an effective viscous pressure or a matter creation pressure. We first focus on the viscous scenario. Under the considerations commented above the pressure of matter sector takes the form
@@
we have used Eq. (4) to write the viscous pressure as $\Pi=\rho_{\Lambda}-\frac{1}{3}\rho_{r}-3H^{2}-2\dot{H}^{2}$. From the degeneracy condition $E(z)=E_{\Lambda}(z)$, one gets the following expression for viscous dark matter.
@@
Therefore, if we insert the result given above in Eq. (19), we get the following expression.
@@
If we consider the conditions given in velten we have $\Omega_{\mathrm{r,0}}=\bar{\Omega}_{\mathrm{r,0}}$ since the density of the radiation fluid is obtained from the temperature of the CMB, in other words, this parameter is not model dependent. Besides, if the correct amount of matter is established correctly from observations then $\Omega_{\mathrm{m,0}}=\Omega_{\mathrm{vm,0}}$, which in turn implies that the dark energy density parameter will be also well determined for any model, then $\Omega_{\mathrm{de,0}}=\Omega_{\Lambda}$. Therefore, Eq. (19) takes the form
@@
where the r.h.s. of Eq. (19) is replaced by the l.h.s. of Eq. (21). Notice that the viscous matter sector can be written in terms of the parameters of the cosmological model (16). It is worth mentioning that the solution for $\Omega_{\mathrm{vm}}(z)$ does not depend on a specific election of $\Pi$, then this solution would be valid to characterize viscous effects obtained from the full theory of the causal Israel-Stewart formalism is and even from the non-linear extension formalism given in nonlinear , which is more adequate to describe an expanding fluid. Since matter creation effects are uniquely characterized by deviations from the equilibrium pressure, Eq. (22) can also describe the created matter scheme, $\Omega_{\mathrm{vm}}(z)\longrightarrow\Omega_{\mathrm{cm}}(z)$, both scenarios will provide a solution for the matter sector in terms of the parameters associated with the decaying dark matter model plus dark energy $(\alpha,\omega_{\mathrm{de}})$ given in Eq. (16). In the context of the matter creation scenario, the Ansatz for the $\Gamma$ term is necessary to determine the form of the creation pressure. However, as can be seen in our description, we can dispense of the explicit form of this term. See for instance Ref. gamma for examples of $\Gamma$.
@@
We would like to highlight the fact that, under the inclusion of viscous effects and the consideration of the degeneracy condition at the background level between both models, the expansion of the universe within the viscous $\Lambda$CDM model (23) behaves exactly as in the decaying dark matter model given in (16). However, the physical interpretation is distinct for each case; as can be seen, the modified $\Lambda$CDM model has a matter sector with a viscous pressure that can be modeled by means of the decaying dark matter model parameters (as we will see below) and the nature of the CC changes into a dynamical behavior, $\Lambda\rightarrow\Lambda(z)$, characterized by the parameter of state $\omega_{\mathrm{de}}$.
In Fig. (1) we show the normalized Hubble parameter (23) and compare it to the $\Lambda$CDM model for the recent evolution of the universe. As can be seen, in the past, the cases are practically indistinguishable for a small value of $\alpha$. However, as the value of $\alpha$ increases, we can observe deviations from $\Lambda$CDM in the past for $z>1$. Some other differences with respect to $\Lambda$CDM arise in the far future ($z=-1$), whereas the Hubble parameter for $\Lambda$CDM tends to the constant value given by, $\sqrt{\Omega_{\Lambda}}$, in our model it vanishes for quintessence and diverges for phantom dark energy, therefore in this phantom scenario the singularity is kicked away, this is known as little rip scenario lr , i.e., the model does not exhibit a true (big rip) singularity at a finite cosmic time, the singularity is only allowed until infinite time has elapsed. See, for instance lr2 , where it is discussed that viscous effects can naturally lead to a little rip cosmic evolution avoiding all problems of a big rip singularity; see also lr3 .
It is worthy to mention that for quintessence dark energy all values of $\alpha$ considered tend to increase the value of the normalized Hubble parameter w.r.t. the $\Lambda$CDM model in the low red-shift region, $10<z<$ 1, this is an interesting characteristic since such behavior is known to alleviate some tensions of $\Lambda$ CDM; see, for instance Ref. vagnozzi where this kind of condition was studied for different dark energy models.
Figure 1: Behavior of the parameter $E_{\Lambda}(z)$. For comparison purposes, we use the values $\Omega_{\mathrm{r,0}}=0$, $\Omega_{\mathrm{m,0}}=0.3$, $\Omega_{\mathrm{de,0}}=0.7$. We focus on the recent evolution of the universe $z\approx 0.6$, since for that redshift value the transition decelerated-accelerated for the cosmic expansion is expected to take place moresco .